OpenStudy (anonymous):

What are the zeroes of f(x) = x2 + 5x + 6?

OpenStudy (irishboy123):

\(x^2 + 5x + 6\) factor it if it is \((x+a)(x+b)\) then you have \(x^2 + (a+b)x + ab\) so that means that \(ab = 6\) and \(a + b = 5\) F.O.I.L.

OpenStudy (imtiaz7):

we'll put f(x)=0 that is: x2 + 5x+ 6=0 we get x=-3 and x=-2 by factorization